### v1

# Using LLMs in Generating Design Rationale for Software Architecture Decisions

Abstract.
Design Rationale (DR) for software architecture decisions refers to the reasoning underlying architectural choices, which provides valuable insights into the different phases of the architecting process throughout software development. However, in practice, DR is often inadequately documented due to a lack of motivation and effort from developers. With the recent advancements in Large Language Models (LLMs), their capabilities in text comprehension, reasoning, and generation may enable the generation and recovery of DR for architecture decisions. In this study, we evaluated the performance of LLMs in generating DR for architecture decisions. First, we collected 50 Stack Overflow (SO) posts, 25 GitHub issues, and 25 GitHub discussions related to architecture decisions to construct a dataset of 100 architecture-related problems. Then, we selected five LLMs to generate DR for the architecture decisions with three prompting strategies, including zero-shot, chain of thought (CoT), and LLM-based agents. With the DR provided by human experts as ground truth, the Precision of LLM-generated DR with the three prompting strategies ranges from 0.267 to 0.278, Recall from 0.627 to 0.715, and F1-score from 0.351 to 0.389. Additionally, 64.45% to 69.42% of the arguments of DR not mentioned by human experts are also helpful, 4.12% to 4.87% of the arguments have uncertain correctness, and 1.59% to 3.24% of the arguments are potentially misleading. Based on the results, we further discussed the pros and cons of the three prompting strategies and the strengths and limitations of the DR generated by LLMs.
Design Rationale, Architecture Decision, Large Language Model, Prompt Engineering, LLM-based Agent
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Software and its engineering Designing softwareccs: General and reference Empirical studies
1. Introduction
The software architecture of a system consists of the structures required to understand the system, including its software elements, their relationships and properties (Bass et al., 2021). When making architecture decisions, software architects need to consider multiple factors like the application domain, architectural styles and patterns, Commercial Off-The-Shelf (COTS) components, and other infrastructure choices required to meet system requirements (Jansen and Bosch, 2005). Design Rationale (DR) explains the reasoning behind architecture decisions, encapsulating the architecture knowledge and thought processes that support the resulting design (Tang et al., 2006). A comprehensive DR can support various development activities, such as change impact analysis or a major redesign. With documented DR, architecture decisions can be more easily revisited and assessed by software architects (Falessi et al., 2013).
Although the importance of DR in architecture design has been widely acknowledged, it is often inadequately recorded and detailed during software development (Tyree and Akerman, 2005; Bosch, 2004). Tang et al. (Tang et al., 2006) conducted a survey to investigate the oversight of DR by developers (e.g., “not aware of” and “no time/budget”) and the challenges of capturing DR (e.g., “no suitable tool”). Besides, extra efforts are required when architects need to capture DR (Capilla et al., 2008). The absence of DR hinders developers from understanding the reasons behind architecture decisions, thus violating the architecture decisions made in the design phase during the implementation and increased maintenance costs (Tang et al., 2007), which reveals a critical gap in the DR capture process during software architecting.
The advancement of Large Language Models (LLMs) offers promising potential to address this gap by assisting in generating DR. LLMs, powered by deep learning technologies, have revolutionized Natural Language Processing (NLP) by demonstrating advanced language understanding capabilities, including syntax, semantics, and pragmatics (Brown et al., 2020). LLMs’ ability to generate diverse content stems from extensive parameters and training on large-scale datasets, enabling sustained performance improvements across a wide range of downstream tasks (Wei et al., 2022a). Previous studies (e.g., (Kabir et al., 2024; Jin et al., 2024; Widjojo and Treude, 2023)) have demonstrated LLMs’ utility in addressing Software Engineering (SE) challenges, highlighting their potential to support developers in various SE tasks. In particular, by leveraging the advanced language comprehension and reasoning capabilities of LLMs, software architects can efficiently use them to assist in generating and recovering DR of architecture decisions. Meanwhile, the accuracy and quality of DR generated by LLMs are yet to be explored. Besides, according to the study by Soliman et al. (Soliman and Keim, 2025), who employed a zero-shot method to query GPT-3.5 regarding the DR of the Hadoop Distributed File System (HDFS) [^1] https://hadoop.apache.org/docs/stable/hadoop-project-dist/hadoop-hdfs/HdfsDesign.html, GPT-3.5 exhibited moderate Recall but relatively low Precision against a predefined ground truth for DR derived from architecture issues and related documents on HDFS. Therefore, how to effectively guide LLMs to ensure the generation of high-quality DR remains a critical challenge that needs further investigation.
Our study aims to explore and validate the feasibility of leveraging LLMs to uncover the DR that underpins the software architecture decisions. By achieving this goal, we seek to enhance the understanding of legacy software systems for developers and stakeholders, thereby ultimately facilitating more effective maintenance and evolution of software systems.
In this paper, we conducted an empirical study to evaluate the DR of architecture decisions generated by LLMs. We first constructed a dataset containing architecture-related problems from software development scenarios. Specifically, the dataset is composed of 100 data points related to architecture decisions, including 50 posts from SO, along with 25 discussions and 25 issues from GitHub. We then chose five widely used LLMs for the evaluation, including both popular open-source and proprietary models: gpt-3.5-turbo, gpt-4-0613, gemini-1.0-pro, llama3-8B, and mistral-7B. To generate DR of architecture decisions, we employed three distinct prompting strategies: zero-shot, chain of thought (CoT), and LLM-based agents. For zero-shot and CoT methods, we designed tailored prompt templates, while for LLM-based agents, we developed a multi-agent system consisting of five LLM-based agents that collaborate to generate DR. Finally, we evaluated the DR generated by the five LLMs and the three prompting strategies using both accuracy metrics for quantitative analysis and IHUM-category for qualitative analysis.
The main contributions of this work are that:
  • We constructed and publicly released a dataset containing 100 architecture-related problems from software development scenarios on SO and GitHub (Zhou et al., 2025).
  • We developed tailored prompt templates for zero-shot and CoT methods, and designed a multi-agent system consisting of five LLM-based agents to generate DR for architecture decisions.
  • We evaluated the accuracy and quality of the DR generated by LLMs through both accuracy metrics and IHUM-category.
  • We discussed the advantages and disadvantages of the three prompting strategies, as well as the strengths and limitations of LLM-generated DR in comparison to that provided by human experts.
Paper Organization: Section 2 introduces the background knowledge relevant to this study. Section 3 presents the Research Questions (RQs) and the research process. Section 4 presents the evaluation results of LLM-generated DR with their interpretation. Section 5 discusses the implications based on the research results. Section 6 clarifies the potential threats to the validity of this study. Section 7 reviews the related work. Finally, Section 8 concludes this work and outlines future research directions.
2. Background
2.1. Definition of Design Rationale
Typically, design rationale in software development refers to the underlying reasons for making design decisions (Moran and Carroll, 2020). However, no standardized definition of DR in software design has been established, resulting in subtle variations in its interpretation across different contexts (Moran and Carroll, 2020). For example, Rogers et al. (Rogers et al., 2015) define that DR consists of the decision problem, alternative solutions, and arguments for or against the solutions. Besides, Zhao et al. (Zhao et al., 2024) defined DR as a combination of a solution and its corresponding supporting or opposing arguments.
We provided the precise interpretation of DR in this study. Our dataset is derived from the posts on SO, as well as the issues and discussions on GitHub. Each data point includes a specific architecture-related problem $\bm{P}$; an architecture decision $\bm{D}$; and its corresponding design rationale $\bm{DR}$. Design rationale ($\bm{DR}$) is composed of a set of arguments $\bm{A}$, where each argument corresponds to one particular perspective for architecture design ($\bm{D}$), such as advantages, disadvantages, and trade-offs. The relationship between architecture problem $\bm{P}$, architecture decision $\bm{D}$, and design rationale $\bm{DR}$ in the dataset is presented in Equation (1) below.
(1) $\{(P,D,DR)\mid DR=\{A_{1},A_{2},\dots,A_{n}\},n\geq 1\}$
2.2. Prompt Engineering of LLMs
Prompt engineering has emerged as an essential technique for enhancing the capabilities of pre-trained LLMs (Sahoo et al., 2024). By designing prompts to guide the output, LLMs can adapt to various tasks and domains without requiring modifications to their parameters (Ratnadira et al., 2024; Deljouyi et al., 2024; Delile et al., 2023). The field of prompt engineering encompasses a wide range of techniques tailored to optimize LLM performance.
In this study, we focused on two prompting methods: zero-shot and CoT. Zero-shot is a foundational prompting technique where the task is directly presented to LLMs without requiring additional examples. This approach enables LLMs to perform tasks by leveraging its pre-training knowledge (Brown et al., 2020). We employed the zero-shot method to evaluate the performance of LLMs in generating DR using the basic prompt. On the other hand, CoT is a technique that guides LLMs through a structured sequence of reasoning steps. Specifically, the template of the CoT method instructs LLMs to generate more thoughtful and deeply reasoned answers by following a series of steps. Wei et al. (Wei et al., 2022b) demonstrated that CoT significantly improves performance in math and commonsense reasoning tasks. Given that DR often involves complex reasoning and trade-offs, we employed CoT in order to help LLMs produce more carefully considered outputs. The templates used for zero-shot and CoT methods are presented in Section 3.
2.3. LLM-based Agents
An intelligent agent is an automated entity capable of perceiving, making decisions, and taking actions (Wooldridge and Jennings, 1995). With the advent of LLMs, researchers and practitioners have increasingly focused on developing LLM-based autonomous agents, leading to the emergence of LLM-based agent frameworks (e.g., BabyAGI (Yohei, 2023) and MetaGPT (Hong et al., 2023)). These agent frameworks typically comprise multiple autonomous intelligent agents, each capable of independently performing tasks by interacting, communicating, and collaborating with other agents. Compared to prompt engineering, multi-agent systems offer more sophisticated agent interactions and coordination mechanisms. By leveraging a variety of tools, LLM-based agent systems can further enhance their capabilities, enabling the systems to collectively address more complex and dynamic challenges. Several LLM-based agent systems have been developed to tackle different tasks in SE, such as code generation (Bairi et al., 2024), code review (Hu et al., 2023), and requirements engineering (Jin et al., 2024). However, there is currently no LLM-based agent systems specifically designed for generating DR for architecture decisions. Considering the complexity of DR, we designed a multi-agent system consisting of five LLM-based agents, building on the MetaGPT framework (Hong et al., 2023). The details of this LLM-based agent system are presented in Section 3.
3. Methodology
In this section, we introduce the methodology employed in our research in terms of research questions (Section 3.1), selection of LLMs (Section 3.2), dataset construction (Section 3.3), design rationale generation (Section 3.4), and evaluation methods (Section 3.5). The overview of the research process is presented in Fig. 1.
Figure 1. Overview of the research process
3.1. Research Questions
Our goal is to uncover and evaluate the capability of LLMs in generating DR for architecture decisions. To this end, we formulated the following three Research Questions (RQs):
RQ1: How effective are LLMs in generating and recovering design rationale for architecture decisions?
Rationale: RQ1 aims to explore the feasibility and capability of using LLMs in generating DR for architecture decisions. DR justifies architecture choices and helps long-term system maintainability by providing context for software engineers to understand architecture decisions (Tang et al., 2007). LLMs can potentially automate the DR generation, helping recover and preserve the reasons behind architecture decisions.
RQ2: Does the CoT technique affect or enhance the capability of LLMs to generate and recover design rationale for architecture decisions?
Rationale: CoT, involving a series of intermediate reasoning steps, significantly enhances the ability of LLMs to perform complex reasoning (Wei et al., 2022b). Since the CoT method enables step-by-step reasoning, it can potentially instruct LLMs to engage in more comprehensive thinking and make more refined trade-offs for architecture decisions. Therefore, we plan to explore the ability of the CoT method to enhance the quality of LLM-generated DR for architecture decisions.
RQ3: Do LLM-based agents affect or enhance the capability of LLMs to generate and recover design rationale for architecture decisions?
Rationale: LLM-based agents have been successfully applied in the field of SE, demonstrating significant effectiveness in supporting and facilitating various SE tasks (Liu et al., 2024b). By collaborating in specific roles, LLM-based agents are capable of performing complicated tasks such as reasoning, gathering information, and reviewing output. However, it remains unclear whether LLM-based agents can effectively generate DR for architecture decisions. Therefore, we designed a multi-agent system to explore whether collaboration among LLM-based agents can further enhance the ability of LLMs to generate DR for architecture decisions.
3.2. Selection of LLMs
As presented in Table 1, we selected five mainstream open-source and closed-source LLMs in this study. These LLMs exhibit high performance in natural language understanding and reasoning, and have been widely applied in many SE tasks (e.g., (Kabir et al., 2023; Kang et al., 2024; de Souza et al., 2023; Zhang et al., 2024)). Besides, the knowledge cut-off date of LLM is also an important factor in our model selection. Since SO posts, GitHub issues and discussions, are publicly available, we collected data generated after the knowledge cut-off dates of all employed LLMs to prevent data leakage issues (Balloccu et al., 2024). Note that selecting an LLM with a more recent cut-off date may result in a smaller dataset, potentially impacting the analysis of experimental results. For instance, if we choose GPT-4o, one of the most advanced models from OpenAI (OpenAI, 2024), only 28 architecture-related posts from SO, which were created after the knowledge cut-off date of GPT-4o, would remain available.
Table 1. Five selected LLMs to generate DR of architecture decisions
Model Description Cut-off date Open-source
gpt-3.5-turbo gpt-3.5-turbo is an LLM from OpenAI’s GPT series. It can understand and generate natural language and code, excelling in tasks like conversation generation, text completion, and question answering. While optimized for chat applications, it also effectively handles non-chat tasks. 2021-09 No
gpt-4-0613 gpt-4-0613 is an advanced multimodal LLM that accepts text and image inputs and produces text outputs. It offers improved accuracy and problem-solving capabilities over previous models, leveraging its extensive knowledge and reasoning skills. gpt-4-0613 is versatile, suitable for chat, text generation, and traditional completion tasks. 2021-09 No
gemini-1.0-pro gemini-1.0-pro is an LLM developed by Google, capable of understanding and generating human language with remarkable proficiency. It can process not only text, but also understand and generate various forms of information such as images and videos. gemini-1.0-pro excels in tasks like multilingual translation, code generation, and question answering. 2023-04 No
llama3-8B llama3-8B is an open-source LLM developed by Meta for efficient natural language processing. It excels in tasks like text understanding, generation, and conversation. With a flexible design, it operates across various platforms and is freely available for researchers and developers to modify. 2023-03 Yes
mistral-7B mistral-7B is an LLM developed by Mistral AI, designed for high-performance natural language processing tasks. It excels in text generation, summarization, and understanding, offering state-of-the-art capabilities for developers and researchers. Additionally, mistral-7B is open-source, allowing for customization and integration into various projects. 2021-08 Yes
3.3. Dataset Construction
As shown in Fig. 2, we constructed a dataset from SO posts, GitHub issues and discussions to facilitate and evaluate the DR generation for practical architecture decisions. To avoid potential data leakage issues, we constructed a new dataset, as the existing datasets may contain data that was used to train the LLMs used in this study as we discussed in Section 3.2. Note that, in our dataset, all posts, issues and discussions were created after the cut-off dates of the five LLMs used in this study. The dataset used in this study has been provided at (Zhou et al., 2025).
Figure 2. Dataset construction regarding architecture decisions
3.3.1. Data Collection
For data collection, we used a keyword-based method to gather data from SO and GitHub. Specifically, we selected the general terms “architect*” (including “architecture”, “architecting”, and “architectural”) to search posts, issues and discussions. Such selection was also used in a recent study to retrieve architecture related SO posts (de Dieu et al., 2023), and it indicates that these terms are effective in capturing more relevant data compared to other general keywords like “design*”. For SO posts, we utilized an SQL query via Stack Exchange Data Explorer interface (Explorer, nd) to collect posts with at least one answer from SO. For GitHub issues and discussions, we collected issues marked as “closed” with the reason specified as “completed”, and discussions marked as “answered”. Note that, we did not limit our search solely to the tags of SO posts, GitHub issues and discussions, as using tags alone may omit some important data related to software architecture topics (Barua et al., 2014). Therefore, we considered both the body and title of posts, issues and discussions during our data search.
Additionally, we limited the retrieval of the posts, issues and discussions to dates after the knowledge cut-off dates of all LLMs. As presented in Table 1, gemini-1.0-pro has the latest cut-off date among the five LLMs we selected, that is, April 2023. Hence, we restrict the data retrieved from SO and GitHub to content after May 2023. We conducted data collection on June 18, 2024, gathering data from SO and GitHub that contained “architect*” in the content body, tags, or title, starting from May 1, 2023. In total, 692 SO posts, 8,050 issues, and 2,546 discussions were retrieved.
3.3.2. Data Review and Labeling
The purpose of the data review and labeling process is to remove data entries that are not applicable to this study. The four criteria for data labeling are as follows:
  • For SO posts, they must include an answer that was either accepted by the question creator or received upvotes from the SO community. This criterion ensures that the architecture decision extracted from the post is sufficiently reasonable to address the current architecture problem.
  • For GitHub issues and discussions, they must come from repositories with more than 200 stars and must not stem from private projects (e.g., coding exercises). This criterion ensures that the architecture decisions come from publicly available projects with a certain level of popularity, thereby reducing the likelihood that the extracted architecture decisions originate from toy projects.
  • The posts, issues, and discussions must be relevant to architecture decisions. Although we collected data that contain “architect*” in tags, titles, or bodies, not all of these data entries are related to software architecture. For example, in certain SO posts, “architecture” refers to “cpu-architecture”, which is focused on hardware rather than software architecture design.
  • The posts, issues, and discussions must explicitly include an “architecture-related problem”, and the corresponding “architecture decision” and “design rationale”. This criterion ensures that the subsequent data extraction process can retrieve the necessary information from these data entries.
The formal data labeling was conducted by the first author. Prior to this, the first and fourth authors randomly selected 25 SO posts, 25 GitHub issues, and 25 GitHub discussions for a pilot labeling process to minimize personal bias and ensure adherence to the data labeling criteria. The inter-rater reliability between the two authors was measured by the Cohen’s Kappa coefficient (Cohen, 1960), yielding values of 0.834. This result indicates an almost perfect agreement between the two authors (Landis and Koch, 1977). The second and third authors were then involved to resolve discrepancies between the pilot results of the first and fourth authors and reach a consensus on the four criteria for data labeling. Subsequently, the first author conducted the formal data labeling according to the four criteria. For SO posts, the first author reviewed and labeled all 692 posts, and identified 50 architecture-related posts for our research. Given the large number of collected GitHub issues and discussions, and to maintain a balanced data volume between SO and GitHub sources, we selected the top 25 issues and top 25 discussions that meet the criteria, sorted by “best match”. Consequently, our dataset comprises 100 entries from SO and GitHub. The data labeling results were reviewed multiple times by the first, second, third, and fourth authors. Discrepancies were addressed through discussions to reach a consensus. Specifically, the first author presented the reasons for the inclusion or exclusion of certain data. The four authors then analyzed and discussed the reasons based on the criteria for data labeling, ultimately reaching an agreement. The pilot and formal data labeling results are recorded in our dataset (Zhou et al., 2025).
3.3.3. Data Extraction
We established a set of data items for data extraction, as presented in Table 2. Data items D1 and D2 are provided as part of the prompt for LLMs and LLM-based agents to generate DR. D3 serves as a reference for evaluation in results analysis, as further elaborated in Section 3.5. The data extraction criteria are as follows:
  • For Architecture Problem (D1), we extract the context related to architecture problems from the SO posts, GitHub issues and discussions as comprehensively as possible.
  • For Architecture Decision (D2), we define different criteria for each data source. For SO, we extract architecture decisions accepted by creators of the posts or received upvotes from the SO community. When multiple architecture decisions satisfy this criterion, we prioritize those accepted by creators; if none are accepted, we choose the one with the highest number of upvotes. For GitHub issues, we extract architecture decisions that have been agreed upon in the issue conversations. For GitHub discussions, we choose architecture decisions that have been marked as the “Answer”.
  • For Design Rationale (D3), we extract the information of the DR related to architecture decisions as comprehensively as possible, including advantages, disadvantages, and relevant trade-offs of architecture decisions. This information forms each argument that constitutes the “Design Rationale”.
The first and fourth authors randomly selected 5 SO posts, 5 GitHub issues, and 5 GitHub discussions to conducted a pilot data extraction. In case of any discrepancies, the second and third authors were involved in reaching a consensus. The results show that the three data items in Table 2 can be extracted from our dataset. Subsequently, the first author conducted the formal data extraction from the labeled dataset. Each extracted data item was reviewed multiple times by the first, second, third, and fourth authors to ensure accuracy. Any inconsistencies between the four authors were discussed according to the criteria of data extraction to achieve a consensus. The data extraction results are recorded in our dataset (Zhou et al., 2025).
Table 2. Extracted data items and their descriptions
# Data Item Description Example (Issue #1256)
D1 Architecture Problem The key point(s) of the architecture problem in SO posts, GitHub issues, or GitHub discussions If you wish to access Console, Router, or Controller via the internet, you must open ports to the public internet and assign a public DNS record. This opens Ziti services to DDoS.
D2 Architecture Decision The key point(s) of the architecture decision in SO posts, GitHub issues, or GitHub discussions I propose implementing UDP hole-punching on the Controller and Router and Tunneler so they can see each other without opening the Controller and Router ports to the Internet. And since the Tunneler can now see the Controller and Router, you can also generate JWT locally on the Tunneller/Client.
D3 Design Rationale The key point(s) of design rationale that consists of the arguments for the architecture decision in SO posts, GitHub issues, or GitHub discussions UDP hole-punching in products such as ZeroTier and Nebula that may be useful for OpenZiti: Argument_1: The supernode is a separate stand-alone service that does not influence sensitive services, such as the controller/router that does the main work and handles and enforces zero-trust, routing, and handle management. So, if the supernode is hacked, the other components will not be compromised. Argument_2: There are many supernodes across the Internet, so it is hard to DDoS all of them simultaneously. Argument_3: Even if all supernodes are DDoSed and down, the peers that already established connection do not care about the supernode and continue working without an issue. Argument_4: Only one port is opened, and you must be concerned about being open on the public Internet to one service that is not viral and acceptable if not working for hours or days.
3.4. Design Rationale Generation
To answer the three RQs (see Section 3.1), we designed prompts for LLMs and LLM-based agents to generate DR based on the data items defined in Table 2. The DR generated with three prompting strategies is recorded in the dataset (Zhou et al., 2025), which are further discussed in Section 4.
3.4.1. Zero-Shot (RQ1)
Zero-shot refers to the capability of a model to understand and complete a task without being specifically trained for it, relying on its generalizability and existing knowledge. Therefore, we only provided “Architecture Problem” and its corresponding “Architecture Decision” in prompts without any example for LLMs to generate DR. To formulate effective templates for the prompts, we conducted a pilot experiment on 10 randomly selected samples from our dataset. During the pilot experiment, we tried several templates. For example, in one template, we added extra information asking LLMs to act as an “architecture decision expert”, detailing their specialties and responsibilities. In another template, we used a shorter instruction simply requesting LLMs to generate DR. By manually experimenting with different prompts on a subset of samples, we identified the most effective template. As shown in Fig. 3, this template helps LLMs understand architecture problems and generate desirable outputs, and its effectiveness was further validated in formal experiments.
Prompt (zero-shot):
I will provide an architecture-related problem and the corresponding architecture decision.
Architecture Problem: {architecture_problem} Architecture Decision: {architecture_decision} Please provide the design rationale of the architecture decision in light of the architecture problem.
Figure 3. A template of the zero-shot method to generate DR
3.4.2. CoT (RQ2)
CoT allows LLMs to progressively demonstrate their thought process before arriving at a final answer, enhancing their reasoning ability and the transparency of their decisions (Wei et al., 2022b). Therefore, in addition to “Architecture Problem” and “Architecture Decision”, we also provided LLMs with a framework for analyzing the rationale in prompts. We designed several versions of the framework for LLMs to analyze the rationale and conducted pilot experiments on 10 randomly selected samples from our dataset. According to the results of the pilot experiments, we identified the most efficient CoT framework with four steps shown in Fig. 4. In Step 1, we ask the LLMs to understand architecture decisions based on the context of architecture problems. In Step 2 and Step 3, we ask the LLMs to analyze the advantages and disadvantages of the architecture decisions, respectively. Finally, in Step 4, we ask the LLMs to present the DR based on the trade-off of the advantages and disadvantages.
It is important to note that we did not provide any other examples in prompts for LLMs. This is because the main purpose of providing additional examples in CoT is to help the model understand the required reasoning process (Wei et al., 2022b). However, the results of the pilot experiments demonstrate that LLMs can generate DR step by step according to the designed CoT framework without any examples, which was further confirmed in the formal experiments. In addition, introducing examples in prompts significantly increases the length of input text, which raises the difficulty for LLMs to understand architecture problems and leads to their generated DR being potentially influenced by specific examples.
Prompt (CoT):
I will provide an architecture-related problem and the corresponding architecture decision.
Architecture Problem: {architecture_problem} Architecture Decision: {architecture_decision} Please provide the design rationale of the architecture decision in light of the architecture problem. Let’s think step by step: Step 1: Understand the architecture decision based on the provided architecture problem. Step 2: Analyze the advantages of choosing this architecture decision. Step 3: Analyze the disadvantages of choosing this architecture decision. Step 4: Considering the advantages and disadvantages with the trade-offs between them, please provide a detailed design rationale for choosing this architecture decision.
Figure 4. A template of the CoT method to generate DR
3.4.3. LLM-based Agents (RQ3)
To answer RQ3 in Section 3.1, we designed five LLM-based agents (i.e., Aspect_Identifier, Information_Collector, Aspect_Analyst, Aspect_Reviewer, and Trade-off_Analyst) by utilizing the MetaGPT framework (Hong et al., 2023). MetaGPT utilizes an assembly line approach to assign different roles to various agents, effectively breaking down complex tasks into subtasks that require collaboration among multiple agents. According to the study of Liu et al. (Liu et al., 2024a), the design pattern of MetaGPT is role-based cooperation, which has been employed in many SE studies (e.g., (Jin et al., 2024; Ratnadira et al., 2024)). Currently, no research has reached a widespread consensus on the role delineation for generating DR. Thus, we approached a process of generating DR from the perspective of architects (Jansen et al., 2008; Shahbazian et al., 2018), breaking it down into a set of subtasks assigned to five distinct agents.
The process for generating DR using LLM-based agents is the following: (1) The architecture problems and the corresponding architecture decisions are provided to the agent system. (2) Aspect_Identifier agent identifies no more than six relevant aspects for analysis. (3) Information_Collector agent gathers relevant background information for each aspect. (4) The aspects and their background knowledge are provided to Aspect_Analyst agent for analysis. (5) The analysis results from Aspect_Analyst agent are reviewed and modified by Aspect_Reviewer agent. (6) Trade-off_Analyst agent compiles all analysis results of the aspects and generates the final DR. (7) Eventually, the final DR is provided as the output of the agent system. The overview of collaboration between the five agents is presented in Fig. 5.
Note: Steps ①-⑦ represent the key steps in the DR generation process based on LLM-based agents. Steps a) - f) represent the specific procedures employed by Information_Collector agent to gather background knowledge for Aspect X.
Figure 5. Overview of collaboration among LLM-based agents to generate DR
The introduction to the five LLM-based agents is as follows:
  • Aspect_Identifier agent is responsible for identifying the most relevant and important aspects related to certain architecture decisions. Specifically, Aspect_Identifier agent receives the architecture problem and the corresponding architecture decision, and then provides up to six of the most relevant aspects, along with their definitions. We set this upper limit at six aspects according to the results of the pilot experiments on 10 samples from our dataset. The results indicated that without such a restriction, LLMs tend to enumerate too many aspects, which can make it difficult to maintain a clear focus in the subsequent analysis of DR. Additionally, this excessive enumeration significantly increases the costs for LLM-based agents. As a result, we found that limiting the aspects to six is considered a reasonable compromise, ensuring both depth of analysis and a concentrated focus. Then, each defined aspect is transmitted to Information_Collector agent.
  • Information_Collector agent is in charge of gathering relevant information through search engines. We initially intended to enhance the information retrieval capabilities of Information_Collector agent by using Retrieval-Augmented Generation (RAG) (Gao et al., 2024) technique with a unified database. However, the diversity of architecture problems prevented us from providing a single database that could adequately address all issues. Consequently, we chose to dynamically retrieve relevant knowledge from search engines based on specific problems. Specifically, Information_Collector agent receives architecture problems, the corresponding architecture decisions, and aspects predefined by Aspect_Identifier agent. Based on each aspect, Information_Collector generates the three most relevant keywords for search engine queries. We chose DuckDuckGo as our search engine as it offers search results without personalized interference and provides a high level of information reliability (Hannak et al., 2013). For each keyword, Information_Collector agent retrieves the top 20 URLs returned by the search engine. Therefore, for each aspect, a total of 60 relevant URLs are collected (i.e., 3 $\times$ 20 URLs). To prevent the agent from using the original SO posts, GitHub issues, or GitHub discussions as reference information (i.e., data leakage issues), we discard any collected background knowledge whose URLs appear in the dataset and replace them with alternative sources. Then, the agent stores the webpage content from these URLs in Facebook AI Similarity Search (FAISS), a library designed for fast nearest-neighbor search and dense vector clustering (Douze et al., 2024). By using the similarity search function of FAISS, Information_Collector agent selects the 10 most relevant pieces of information from all the collected webpages that are most similar to the given architecture problems and architecture decisions. The 10 pieces of information, are then transmitted to Aspect_Analyst agent as background knowledge along with the aspects and their definitions identified by Aspect_Identifier agent. This collected information has not undergone rigorous manual checking, and as a result, it may contain irrelevant or inaccurate details. We addressed this issue through Aspect_Analyst and Aspect_Reviewer agents.
  • Aspect_Analyst agent is tasked with analyzing different aspects of architecture decisions. Specifically, Aspect_Analyst agent inputs are architecture problems, the corresponding architecture decisions, aspects that need to be analyzed, and the background knowledge collected by Information_Collector agent. Then, Aspect_Analyst agent provides a detailed analysis of the chosen architecture decisions in each aspect. Each analysis for different aspects is transmitted to Aspect_Reviewer agent. Additionally, since the background knowledge may contain inaccurate information, we alerted Aspect_Analyst agent to this risk and instructed it to verify and review the information before using it.
  • Aspect_Reviewer agent is accountable for reviewing the analysis results of different aspects from Aspect_Analyst agent. Specifically, Aspect_Reviewer agent receives architecture problems, the corresponding architecture decisions, and relevant aspects along with their analysis. Note that, we did not provide the background knowledge to Aspect_Reviewer agent. This decision is based on our pilot experiments, which indicate that introducing background knowledge could expose Aspect_Reviewer agent to potential biases, hindering the ability of LLMs to maintain an independent and objective evaluation. If Aspect_Reviewer agent considers the analysis to be reasonable, the analysis results will be accepted. Otherwise, Aspect_Reviewer agent would identify the deficiencies in the analysis results and propose necessary modifications to the analysis results. The analyses that either pass the review or are modified accordingly will be forwarded to Trade-off Analyst agent for final evaluation.
  • Trade-off_Analyst agent needs to provide a final DR for the chosen architecture decisions based on all previous analysis results. Specifically, Trade-off_Analyst agent receives architecture problems, the corresponding architecture decisions, and the reviewed analysis results provided by Aspect_Reviewer agent. Then, Trade-off_Analyst agent analyzes the trade-offs to evaluate the architecture decisions and provides a final DR explaining why this specific design decision was made.
Liu et al. (Liu et al., 2024a) identified that the main drawback of role-based collaboration in multi-agent design involves increased communication overhead due to collaboration between agents, and the possibility of varying prices for agent services with different roles. In our pilot experiments, we found that making the interaction process of agents overly complex did not greatly improve the quality of the generated DR. Therefore, we referenced the single-path plan generator design pattern (Liu et al., 2024a), making each agent have only one subsequent step after completing its task, which is to submit the necessary information from the task execution results to the next agent.
3.5. Evaluation Methods
We employed both accuracy metrics and IHUM-category to evaluate the generated DR. Accuracy metrics, including Precision, Recall, and F1-score, are employed for quantitative analysis. IHUM-category, a classification we defined in this study, is used for the qualitative analysis of DR, which is detailed in Section 3.5.2. The pilot evaluation was conducted by the first and fourth authors. They randomly selected 15 DR generated by LLMs to conduct a pilot evaluation: five generated with zero-shot, five with CoT, and five with LLM-based agents. The inter-rater reliability between the two authors was measured by the Cohen’s Kappa coefficient (Cohen, 1960). The Cohen’s Kappa value for the accuracy metrics is 0.863, whereas for IHUM-category is 0.738, indicating almost perfect and substantial agreement between the two authors, respectively. Then, the second and third authors joined the discussion with the first and fourth authors to resolve discrepancies and reach a consensus on the evaluation criteria. Subsequently, the first author conducted a formal evaluation on all generated DR. The evaluation results were thoroughly reviewed by the four authors (the first to fourth authors), with discrepancies resolved through discussion to reach a consensus. It is important to note that we did not use text similarity and semantic similarity metrics such as BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), and METEOR (Banerjee and Lavie, 2005), to evaluate the LLM-generated DR. Based on the pilot experiments conducted on 10 randomly selected DR generated by LLMs, we observed a significant difference between the text provided by human experts and that generated by LLMs. Simply applying text and semantic similarity metrics yields very low values with all the three prompting strategies, making it challenging to draw valid experimental conclusions.
3.5.1. Accuracy Metrics
For accuracy metrics, we used Precision, Recall, and F1-score to evaluate the alignment of DR generated by LLMs with those provided by human experts. The three metrics are commonly used to measure the quality of DR (e.g., (Zhao et al., 2024; Soliman and Keim, 2025)). As mentioned in Section 2, DR comprises a set of arguments. We compared the arguments provided in the DR generated by LLMs with those extracted from human experts during the data extraction phase. We then labeled the arguments into true positive (TP), false positive (FP), and false negative (FN). To be specific, we set the following criteria for labeling arguments:
  • TP: The argument points present the DR provided by both LLMs and human experts.
  • FP: The argument points that are present exclusively in the DR generated by LLMs and do not align with the ground truth provided by human experts.
  • FN: The argument points that are missing from the DR generated by LLMs but are present in the ground truth provided by human experts.
  • TN: The argument points that are absent from both the LLM-generated DR and the ground truth provided by human experts. Such argument points are not included in our evaluation, as they do not exist.
Equations  (2),  (3) and  (4) present three accuracy metrics, Precision, Recall, and F1-score. For a given data point $(P_{i},D_{i},DR_{i})$, $TP_{i}$, $FP_{i}$, and $FN_{i}$ are the number of arguments classified as true positive, false positive, and false negative in the DR. As the number of arguments from human experts is typically small (usually less than three), slight variations in the classification of arguments as $TP$, $FP$, and $FN$ can cause significant fluctuations in the metrics. For instance, when two arguments are provided by an expert, an increase in $TP_{i}$ from one to two results in a sharp rise in Recall from 50% to 100%. Therefore, these accuracy metrics are derived by calculating the weighted average based on the number of arguments provided by human experts for 100 data points.
(2) $\text{Precision}=\frac{\sum(TP_{i})}{\sum(TP_{i}+FP_{i})}$
(3) $\text{Recall}=\frac{\sum(TP_{i})}{\sum(TP_{i}+FN_{i})}$
[…]
3.5.2. IHUM-Category
We introduced the IHUM-category to qualitatively assess the argument points in DR generated by LLMs. Accuracy metrics are used to compare the similarity between the DR generated by LLMs and that provided by human experts. However, there may be a chance that the responses generated by LLMs do not align with those of human experts (i.e., FN); the arguments still offer valuable insights and provide alternative perspectives. We proposed that the evaluation of LLM-generated DR could follow a methodology akin to the one employed by Mahajan et al. (Mahajan et al., 2020) for assessing responses on SO. Therefore, we introduced the IHUM-category, which is adapted from the metrics in (Mahajan et al., 2020), with adjustments made to effectively assess DR. IHUM-category is used to manually classify all the argument points, allowing for a more comprehensive evaluation of the DR generated by LLMs. The definition of IHUM-category is presented in Table 3.
Table 3. IHUM-category for classifying argument points in LLM-generated DR
Label Definition
Insightful (I) The argument provides correct and valuable perspectives that are most relevant to the architecture decision. Given the difficulty of assessing the relevance of an argument to the architecture decision, we refer to the rationale provided by human experts. Specifically, if the argument generated by LLMs aligns with the argument points raised by human experts, we consider it to be “Insightful”.
Helpful (H) The argument provides a correct perspective that contributes to the architecture decision, but may not be the most critical reason for making this design decision.
Uncertain (U) The argument is difficult to assess for correctness due to vague descriptions or insufficient contextual information.
Misleading (M) The argument provides an incorrect or misleading perspective.
4. Results and Interpretation
In this section, we present the results of the three RQs with their interpretation. Specifically, we provide the accuracy metrics and IHUM-category classification results of the DR generated by the five LLMs (see Section 3.2) using three distinct prompting strategies (see Section 3.4), detailed in Tables 45, and 6, respectively. Table 7 presents the average performance of the five LLMs in generating DR with three prompting strategies. In Tables 4 to 7, the best-performing results among the five LLMs are highlighted using both underlining and bold text. Besides, since the number and proportion of arguments in the “Uncertain” and “Misleading” categories are negative indicators, the smallest values are highlighted in the tables.
4.1. RQ1: How effective are LLMs in generating and recovering design rationale for architecture decisions?
Table 4 presents the accuracy metrics (including Precision, Recall, and F1-score) and IHUM-category classification results of the DR generated by the five LLMs with the zero-shot method.
Table 4. The evaluation results of DR generated with the zero-shot method
gpt-3.5-turbo gpt-4-0613 gemini-1.0-pro llama3-8B mistral-7B Average
Precision 0.265 0.250 0.295 0.286 0.277 0.278
Recall 0.591 0.697 0.652 0.576 0.621 0.627
F1-score 0.366 0.368 0.406 0.382 0.383 0.381
Argument count 441 553 438 399 444 455
Insightful (I) 117 138 129 114 123 124.2
Helpful (H) 290 384 273 246 280 294.6
Uncertain (U) 22 23 22 19 24 22
Misleading (M) 12 8 14 20 17 14.2
I/(I+H+U+M) 26.53% 24.96% 29.45% 28.57% 27.70% 27.44%
H/(I+H+U+M) 65.76% 69.44% 62.33% 61.65% 63.06% 64.45%
U/(I+H+U+M) 4.99% 4.16% 5.02% 4.76% 5.41% 4.87%
M/(I+H+U+M) 2.72% 1.45% 3.20% 5.01% 3.83% 3.24%
4.1.1. For Accuracy Metrics
The average Precision, Recall, and F1-score of the DR generated by the five LLMs with the zero-shot method are 0.278, 0.627, and 0.381, respectively. The results show that approximately one-quarter of the arguments in the DR generated by the LLMs align with those provided by human experts. Additionally, over 60% of the expert-provided arguments are also mentioned in the LLM-generated DR. Among the five LLMs, gemini-1.0-pro exhibits the highest Precision (0.295) and F1-score (0.406), while gpt-4-0613 achieves the highest Recall (0.697). The relatively lower Precision and F1-score are mainly due to the higher number of arguments in the generated DR, which is approximately 2.3 times greater than those provided by human experts.
4.1.2. For IHUM-Category
The average proportion of arguments classified under “Insightful”, “Helpful”, “Uncertain”, and “Misleading” categories are 27.44%, 64.45%, 4.87%, and 3.24%, respectively. It is worth noting that besides the 27.44% arguments that are consistent with human experts (“Insightful”), around 65% are not considered by the human experts, but also offer “Helpful” insights into architectural design. Additionally, 4.87% of the arguments are indeterminate (“Uncertain”) and 3.24% contain potential errors (“Misleading”), representing a relatively small proportion. Among the five LLMs, gemini-1.0-pro generated the highest proportion of “Insightful” arguments (29.45%), while gpt-4-0613 produces the highest proportion of “Helpful” arguments (69.44%), along with the lowest proportions in the “Uncertain” (4.16%) and “Misleading” categories (1.45%).
Key findings of RQ1 With the zero-shot method, the five LLMs are capable of generating DR for the architecture decisions in our dataset. About 28% of the arguments in the DR generated by LLMs align with the insights from human experts, representing 62% of the total arguments the human experts provided. The higher number of arguments in the DR generated by the LLMs compared to those provided by human experts is the reason for its relatively low Precision and F1-score. About 65% of the arguments in the generated DR provide “Helpful” insights that were not considered by human experts. Besides, fewer than 5% of the arguments are “Uncertain” in correctness, while 3.24% are “Misleading”.
4.2. RQ2: Does the CoT technique affect or enhance the capability of LLMs to generate and recover design rationale for architecture decisions?
Table 5 presents the accuracy metrics (including Precision, Recall, and F1-score) and IHUM-category classification results of the DR generated by the five LLMs with the CoT method.
Table 5. The evaluation results of DR generated with the CoT method
gpt-3.5-turbo gpt-4-0613 gemini-1.0-pro llama3-8B mistral-7B Average
Precision 0.280 0.231 0.224 0.235 0.214 0.237
Recall 0.682 0.717 0.657 0.667 0.697 0.684
F1-score 0.396 0.350 0.335 0.347 0.327 0.351
Argument count 483 614 578 562 646 576.6
Insightful (I) 135 142 130 132 138 135.4
Helpful (H) 314 442 400 389 463 401.6
Uncertain (U) 25 23 30 23 30 26.2
Misleading (M) 9 7 18 18 15 13.4
I/(I+H+U+M) 27.95% 23.13% 22.49% 23.49% 21.36% 23.68%
H/(I+H+U+M) 65.01% 71.99% 69.20% 69.22% 71.67% 69.42%
U/(I+H+U+M) 5.18% 3.75% 5.19% 4.09% 4.64% 4.57%
M/(I+H+U+M) 1.86% 1.14% 3.11% 3.20% 3.32% 2.33%
4.2.1. For Accuracy Metrics
The average Precision, Recall, and F1-score of the DR generated by the five LLMs with the CoT method are 0.237, 0.684, and 0.351, respectively. Among the five LLMs, gpt-3.5-turbo demonstrates the highest Precision (0.280) and F1-score (0.396), while gpt-4-0613 attains the highest Recall (0.717).
Except for gpt-3.5-turbo, the performance of the other four LLMs in terms of Precision and F1-score deteriorates with the CoT method, compared to the performance of the zero-shot method. This can be explained by the significantly higher number of arguments generated in the DR when using the CoT template, reaching 5.76 per DR on average, compared to 4.55 per DR with the zero-shot method. The increased number of arguments stems from the CoT template explicitly guiding LLMs to conduct a trade-off analysis after providing both the advantages and disadvantages. In contrast, the DR generated with the zero-shot template is less likely to consider the negative impacts of architecture decisions. Due to the increase in the number of arguments generated with the CoT template, the DR considers more diverse perspectives compared to the zero-shot method, which leads to a 9.1% increase in Recall. Additionally, the decrease in Precision and F1-score is also primarily attributed to the increase in the total number of arguments generated in the DR, which leads to a lower proportion of arguments consistent with those provided by human experts. However, this does not imply a decline in the quality of DR generated using the CoT method, as we cannot expect the LLM-generated DR to consist solely of arguments provided by human experts. Arguments not considered by human experts in the LLM-generated DR can also potentially offer valuable insights, which are further evaluated according to the IHUM category in Section 4.2.2.
4.2.2. For IHUM-Category
The average proportion of arguments classified under “Insightful”, “Helpful”, “Uncertain” and “Misleading” categories are 23.68%, 69.42%, 4.57%, and 2.33%, respectively. Among the five LLMs, gpt-3.5-turbo exhibits a notable increase in the proportion of “Insightful” arguments (27.95%), ranking the highest. Meanwhile, gpt-4-0613 maintains the highest proportion of “Helpful” (71.99%) arguments and has the lowest proportion of “Uncertain” (3.75%) and “Misleading” arguments (1.14%).
Table 7 presents a comparison of the average performance between the zero-shot and CoT methods. Although the proportion of “Insightful” arguments decreases, the absolute number of “Insightful” arguments generated with the CoT method increases, which is also reflected in the higher average Recall. Additionally, compared to the zero-shot method, the number of “Helpful” arguments increases by 36.3%, while the proportion increases by 7.7%. This improvement indicates that, with the CoT method, LLMs not only generate more “Insightful” arguments but also present a broader range of “Helpful” arguments, offering architects a more comprehensive perspective on architecture decisions. However, in certain DR, such as the one generated by gpt-4-0613 for SO #78536441 and mistral-7B for Discussion #9022, we identified an excess of arguments, which are 14 and 10 respectively. Some arguments that are less relevant to the architecture decision may result in the DR appearing overly lengthy and unfocused. Potential issues of the CoT method are further discussed in Section 5.1.
The proportion of “Uncertain” arguments has not significantly decreased. However, the proportion of “Misleading” arguments in the DR has declined. It is noticed that the CoT method enhances reasoning process by outlining the strengths and drawbacks of architecture decisions for conducting a trade-off analysis, thereby mitigating some of the errors observed in the zero-shot approach. For example, SO #77745340 examines the rationale for “running .NET projects targeting different framework versions in separate processes”, with Inter Process Communication (IPC) mechanisms implied. The DR generated by gpt-3.5-turbo using the zero-shot method suggests that “running components in separate processes can utilize system resources more efficiently”, concluding that the whole system performance is enhanced. However, the introduction of IPC actually increases the system’s performance costs due to the need for serialization and deserialization during data transmission between different processes. While the DR generated with the zero-shot method ignores this issue (the cost of IPC), the DR generated by the CoT method considers the issue when analyzing the disadvantages of this architecture, ultimately leading to a more accurate trade-off conclusion.
Key findings of RQ2 Compared to the zero-shot method, the DR generated through the CoT method includes more arguments aligned with human experts, leading to an increase in Recall. However, the higher number of arguments caused a decrease in both Precision and F1-score. Since the explicit specification in the CoT template to analyze both advantages and disadvantages of architecture decisions, the generated DR provides a more comprehensive perspective, including a greater number and proportion of “Helpful” arguments compared to the zero-shot method. By analyzing the pros and cons of architecture decisions and incorporating a trade-off, the CoT method avoids some errors introduced by the zero-shot method, leading to fewer “Misleading” arguments.
4.3. RQ3: Do LLM-based agents affect or enhance the capability of LLMs to generate and recover design rationale for architecture decisions?
Table 6 presents the accuracy metrics (including Precision, Recall, and F1-score) and IHUM-category classification results of the DR generated by LLM-based agents.
Table 6. The evaluation results of DR generated by LLM-based agents
gpt-3.5-turbo gpt-4-0613 gemini-1.0-pro llama3-8B mistral-7B Average
Precision 0.291 0.255 0.269 0.254 0.266 0.267
Recall 0.717 0.747 0.707 0.687 0.717 0.715
F1-score 0.414 0.380 0.389 0.371 0.386 0.389
Argument count 488 580 521 535 533 531.4
Insightful (I) 142 148 140 136 142 141.6
Helpful (H) 318 407 345 365 363 359.6
Uncertain (U) 21 20 26 21 21 21.8
Misleading (M) 7 5 10 13 7 8.4
I/(I+H+U+M) 29.10% 25.52% 26.87% 25.42% 26.64% 26.71%
H/(I+H+U+M) 65.16% 70.17% 66.22% 68.22% 68.11% 67.58%
U/(I+H+U+M) 4.30% 3.45% 4.99% 3.93% 3.94% 4.12%
M/(I+H+U+M) 1.43% 0.86% 1.92% 2.43% 1.31% 1.59%
4.3.1. For Accuracy Metrics
The average Precision, Recall, and F1-score of the DR generated by LLM-based agents are 0.267, 0.715, and 0.389, respectively. Among the five LLM-based agent systems, gpt-3.5-turbo achieves the highest Precision (0.291) and F1-score (0.414), while gpt-4-0613 demonstrates the highest Recall (0.747).
In Table 7, LLM-based agents generate more arguments than zero-shot but fewer than the CoT method, as Aspect_Identifier agent is required to specify up to six most relevant perspectives for architecture analysis. The DR generated by LLM-based agents has a lower average Precision than the zero-shot method but higher than the CoT method, while its F1-score is comparable to the zero-shot method and superior to the CoT method. LLM-based agents show a notable improvement in Recall, with an increase of 19.8% compared to the zero-shot method and 4.5% compared to the CoT method. This result indicates that LLM-based agents can identify and focus on the key factors that align with human experts’ arguments in architecture decisions.
4.3.2. For IHUM-Category
The average proportion of arguments classified under “Insightful”, “Helpful”, “Uncertain” and “Misleading” categories are 26.71%, 67.58%, 4.12%, and 1.59%, respectively. Among the five LLMs, gpt-3.5-turbo exhibits an increase in the proportion of “Insightful” arguments (29.10%), ranking the highest. Meanwhile, gpt-4-0613 maintains the highest proportion of “Helpful” arguments (70.17%) and has the lowest proportion of “Uncertain” (3.45%) and “Misleading” arguments (0.86%).
Compared to zero-shot and CoT, LLM-based agents generate the highest number of “Insightful” arguments in the DR, achieving 1.42 on average per DR. Additionally, we also identified a substantial proportion of “Helpful” arguments, approximately 68%, which is higher than the zero-shot method but slightly lower than the CoT method.
The proportion of “Uncertain” and “Misleading” arguments in the DR generated by LLM-based agents has decreased compared to other prompting strategies. The decline in “Uncertain” arguments is relatively marginal, while the reduction in “Misleading” arguments is more significant, with the proportion decreasing to 1.59%. According to the output of each agent, the reduction in “Misleading” arguments is attributed to Aspect_Reviewer and Trade-off_Analyst agents. Aspect_Reviewer agent plays a critical role in identifying and rectifying errors within the arguments. For example, in the DR generated by gpt-3.5-turbo for Discussion #787, Aspect_Reviewer agent highlights that the performance analysis by Aspect_Analyst lacks sufficient rigor, emphasizing that factors such as concurrency, resource management, and scalability are also crucial in evaluating whether the architecture decision would lead to performance improvements. This analysis of Aspect_Reviewer agent effectively prevents the inclusion of vague and inadequately substantiated conclusions of “enhancing system efficiency” as a reference in the final trade-off analysis. Therefore, in the final DR, the performance discussion focuses on potential concurrency issues in a multi-threaded environment and the need for effective resource management, rather than claiming a definitive performance improvement. Trade-off_Analyst agent is required to conduct a trade-off analysis according to the analysis from different aspects and provide the final DR. We observed that the DR generated by the multi-agent system based on five LLMs incorporates a consideration of both positive and negative impacts, which is similar to the DR generated with the CoT template.
Key findings of RQ3 Compared to the zero-shot and CoT methods, the DR generated by LLM-based agents includes the most arguments align with human experts, exhibiting the highest Recall. The Precision is slightly lower than that of the zero-shot method due to the increased arguments number, while the F1-score is comparable to the zero-shot method. The increase in “Insightful” arguments and the decrease in “Helpful” arguments ensure that the DR generated by LLM-based agents provides a comprehensive perspective on the architectural decision while minimizing the inclusion of low-relevance arguments. The DR generated by LLM-based agents includes the fewest “Uncertain” and “Misleading” arguments, primarily due to the modifications made by Aspect_Reviewer agent and the analysis provided by Trade-off_Analyst agent.
Table 7. Comparison of evaluation results from the three prompting strategies
zero-shot CoT LLM-based Agents
Precision 0.278 0.237 0.267
Recall 0.627 0.684 0.715
F1-score 0.381 0.351 0.389
Argument Count 455 576.6 531.4
I/(I+H+U+M) 27.44% 23.68% 26.71%
H/(I+H+U+M) 64.45% 69.42% 67.58%
U/(I+H+U+M) 4.87% 4.57% 4.12%
M/(I+H+U+M) 3.24% 2.33% 1.59%
5. Discussion
5.1. Pros & Cons of Three Prompting Strategies
[…]
Zero-shot can generate effective DR with single-round dialogue, however, its key limitation lies in the lack of comprehensive analysis and the inclusion of some irrelevant content.
Zero-shot is a basic prompting strategy that provides LLMs with the context of architecture decisions and basic instructions to generate DR, requiring only one single LLM invocation. According to the average performance presented in Table 7, the DR generated with zero-shot method demonstrates the highest Precision (0.278) and an F1-score (0.381) comparable to that of LLM-based agents. For other metrics, the zero-shot approach also shows no significant gap compared to other methods. The performance of the zero-shot method indicates that LLMs are capable of generating DR for architecture decisions without additional training.
However, the DR generated through zero-shot rarely identifies the potential issues of architecture decisions for analysis. For the five LLMs, only 16% of the DR generated by the zero-shot method includes an analysis of architectural weaknesses. Therefore, the arguments provided in the DR generated by zero-shot may lack comprehensive analysis. For instance, the DR generated for SO #77745340 by gpt-3.5-turbo incorrectly judges the performance of the architecture decision due to overlooked IPC overhead, as mentioned in Section 4.2.2. The lack of analysis of potential architectural weaknesses may result in skewed conclusions in arguments, which contributes to the highest proportion of “Misleading” arguments in the DR generated with the zero-shot method. In addition, the prompt template for the zero-shot method does not specify any requirements for the output. As a result, the analysis of DR generated by the zero-shot method is unpredictable and may occasionally include content unrelated to the DR. For example, the DR generated by gemini-1.0-pro for Issue #45 also suggests other architecture decisions to mitigate “thundering-herd scenarios”, where multiple clients or processes simultaneously compete for a shared resource. Such information in the output is unrelated to the DR and may not align with the expectations of software architects regarding the generated content.
[…]
CoT enhances the comprehensiveness of DR by incorporating trade-off analysis compared to the zero-shot method, however, it lacks in-depth analysis for individual arguments.
The CoT method instructs LLMs to generate DR step by step according to the provided template. All five LLMs correctly follow the instructions in the CoT template during DR generation, outlining the potential advantages and disadvantages of architecture decisions for the trade-off analysis in the final step. Compared to the zero-shot method, the CoT method enhances the reasoning capabilities in DR generation, enabling more thorough analysis during the trade-off process and reducing the proportion of “Misleading” arguments compared to the zero-shot method, as explained in Section 4.2.2.
However, some argument points in the DR generated using the CoT templates lacks the analytical depth. According to Table 7, the number of argument points generated by CoT is 26.7% higher than that generated by the zero-shot method and 8.5% higher than that generated by LLM-based agents. Although these arguments provide diverse perspectives on understanding architecture decisions, some of them are merely superficial statements without in-depth analysis and reasoning. For example, in the generated DR for Issue #2378 by gpt-3.5-turbo with the CoT method, the discussions on “increased code complexity” just briefly mention the difficulty of implementing the architecture without a detailed analysis of the specific challenges or potential implications of this added complexity. Similarly, the generated DR also does not provide specific analysis when considering “future reusability”. Therefore, software architects could struggle to derive valuable insights from such argument points due to the lack of in-depth analysis.
[…]
LLM-based agents use multi-agent collaboration for information gathering, reasoning, and verification, however, excessive interactions between agents increase conversation rounds with LLMs, and certain agent functionalities remain to be further optimized.
According to Fig. 5, the reasoning process of LLM-based agents involves two steps: Aspect_Analyst agent analyzes each aspect related to architecture decisions, while Trade-off_Analyst agent provides the final DR based on the analysis of these aspects. By analyzing each architecture-related aspect individually, Aspect_Analyst can provide a more detailed analysis for every aspect, mitigating the issue in CoT where arguments are outlined without sufficient analysis. Trade-off_Analyst agent is tasked with performing a trade-off analysis based on the evaluation of different aspects of architecture decisions, which have been reviewed by Aspect_Reviewer agent beforehand. Therefore, the final DR generated by Trade-off_Analyst agent also includes an analysis of the advantages and disadvantages of architecture decisions, similar to the CoT method, ensuring the comprehensiveness of the generated DR. By leveraging these two layers of reasoning (i.e., Aspect_Reviewer and Trade-off_Analyst agents), LLM-based agents generate DR that better captures key concerns of human experts in architecture decisions than CoT. According to Table 7, the DR generated by LLM-based agents exhibits the highest Recall (0.715) among the three prompting strategies, which indicates that it includes most arguments that align with those of human experts. Besides, the review process executed by Aspect_Reviewer, which checks and modifies the analysis results of each aspect by Aspect_Analyst agent, leads to a decreased proportion of “Misleading” arguments in the generated DR as explained in Section 4.3.2.
However, one of the issues of LLM-based agents in generating DR is the increased conversation rounds with LLMs. The generation of one DR may require up to 20 rounds of conversation with the LLMs, which is significantly greater than required by the zero-shot and CoT methods. In addition, some functions of the agents can be further optimized. For instance, it is challenging to assess the value of external knowledge provided by Information_Collector agent. Although certain DR, such as the one for Issue #1517 generated by mistral-7B, explicitly integrates “recent” findings to support the arguments, the impact of external knowledge is difficult to discern in most cases. Therefore, Aspect_Analyst could incorporate functionality for tracking external knowledge in order to clearly identify which arguments are derived from what external knowledge. Furthermore, compared to human experts, the arguments in the LLM-generated DR typically lack concrete code examples. Therefore, equipping Information_Collector with the ability to automatically retrieve open-source project code or code from Q&A sites relevant to the architecture problem, and enabling Aspect_Analyst to generate context-based code examples, would help LLM-based agents to provide DR closer to practical development scenarios.
5.2. Strengths and Limitations of LLM-Generated DR
[…]
Strength 1: LLMs are capable of identifying the key factors of DR considered by human experts. During the evaluation process, the arguments of DR provided by human experts are regarded as ground truth to quantitatively evaluate the accuracy metrics of LLM-generated DR. Table 7 shows that LLMs exhibit an average Recall higher than 0.6 with all three prompting strategies, in which LLM-based agents even exceed 0.7. This result reveals that most of the key points considered by human experts when making architecture decisions are identified in LLM-generated DR. To enable LLMs to perform multidimensional analysis from different aspects for each architecture problem, we did not explicitly specify in the prompt which perspectives should be focused on. Therefore, the Recall results suggest that LLMs have the potential to reason the key factors of DR based on the specific context of architecture problems and their corresponding architecture decisions.
[…]
Strength 2: LLMs tend to provide more arguments from diverse perspectives in DR than human experts.
The number of arguments in the LLM-generated DR is significantly higher than that provided by human experts. For each architecture decision, human experts provide an average of 1.98 arguments in DR, while LLMs generate an average of 4.55, 5.77, and 5.31 arguments using the three prompting strategies (i.e., zero-shot, CoT, LLM-based agents), respectively. Despite covering most of the arguments provided by human experts, the greater number of arguments in LLM-generated DR leads to a Precision below 0.3 and an F1-score below 0.4. The result suggests that LLMs typically analyze a broader range of aspects in DR for architecture decisions compared to human experts. Human experts may prioritize the feasibility of architecture implementation and its alignment with the project requirements, potentially overlooking other critical considerations. In SO #77579165, the expert suggested “creating a concrete repository class and mocking it for testing in Riverpod and Flutter”, simplifying implementation by synchronizing two files for efficiency and clear responsibility separation. In contrast, the DR generated by llama3-8B expands on this by further emphasizing the separation of concerns, testability, and reusability. Thus, an additional value of LLM-generated DR is the potential ability to complement the DR provided by human experts and help architects to understand architecture decisions from diverse perspectives.
[…]
Limitation 1: DR generated by LLMs may lack the practical project experience and context-specific details that human experts provide.
While LLMs can generate DR from perspectives aligned with those of human experts, they often lack the depth of practical experience and contextual understanding, which are essential characteristics of human expertise. For example, in Issue #65, “Debian” was chosen as the operating system for the container image as an architecture decision. This decision was explained by human experts who not only emphasized Debian’s advantages, but also considered the time loss associated with code refactoring incurred due to migrations from other Linux distributions (e.g., Amazon Linux 2, Oracle Linux, and Universal Base Image) to Debian-based systems. This practice-based analysis is more persuasive for supporting the architecture decision, whereas the DR generated by LLMs typically lacks such experience-driven arguments. In Discussions #442, while the DR generated by LLMs also addressed the complexity of implementing the architecture at the code level, human experts were able to explicitly identify specific potential technical barriers according to the architecture context. Therefore, enabling LLMs to exhibit the experience-driven insights and contextual sensitivity of human experts when analyzing arguments is crucial to improving the quality of the generated DR.
[…]
Limitation 2: The proportion of “Uncertain” arguments in the DR generated by LLMs is difficult to reduce, primarily due to constraints in the provided context and insufficient domain-specific details.
The proportion of “Uncertain” arguments is 4.87%, 4.57%, and 4.12% with the three prompting strategies, respectively. Compared to the zero-shot method, neither the CoT method nor LLM-based agents significantly reduce the proportion of “Uncertain” arguments. A potential reason is the insufficient contextual information related to the architecture. Although we provided as much detailed contextual information as possible collected from the data extraction phase, the available information for LLMs to generate DR remains limited. For instance, regarding architecture problems in GitHub projects, LLMs lack critical insights that project developers possess, such as the previous versions and specific code contexts. Moreover, the challenges that architecture design faces in real-world scenarios remain unknown to LLMs. As a result, the “Uncertain” arguments proposed by LLMs require validation in practical production environments. However, “Uncertain” arguments may also offer valuable insights, encouraging architects to further analyze and consider them in their projects context.
6. Threats to Validity
In this section, we discussed the threats to validity following the guidelines by Runeson and Höst (Runeson and Höst, 2009). It should be mentioned that internal validity is excluded as we did not explore the causal relationships between variables and outcomes.
Construct Validity. The main threat to the construct validity of this study is the evaluation methods we employed to assess the quality of LLM-generated DR. Although we selected architecture decisions with DR provided by human experts in our dataset, it is important to note that there is typically no universally definitive DR for a given architecture decision. Therefore, we employed two complementary methods for evaluating the generated DR: accuracy metrics and the IHUM-category. For accuracy metrics, we compared the arguments in the generated DR with those provided by human experts to calculate the Precison, Recall and F1-score value. Since the DR provided by human experts is used to explain architecture decisions, the values of the accuracy metrics can reflect the capability of LLMs to recover the DR for architecture decisions. However, accuracy metrics considered the DR provided by human experts as the ground truth, overlooking the arguments in LLM-generated DR that may differ from those of human experts. Therefore, we proposed the IHUM-category to qualitatively assess the arguments in the generated DR, categorizing them into those that align with the human experts’ arguments (“Insightful”), those that differ but offer valid perspectives (“Helpful”), those that are ambiguous (“Uncertain”), and those contain erroneous information (“Misleading”). We evaluated the quality of LLM-generated DR by using both methods simultaneously, aiming to assess generated DR as comprehensively as possible.
External Validity. The primary threat to the external validity is the selection of data sources. In our dataset, we chose the posts from SO, as well as the issues and discussions from GitHub to maximize the diversity of data sources. SO is a popular Q&A community where developers engage in discussions, including a wide range of architecture-related issues (de Dieu et al., 2023). GitHub Issues are used for tracking bugs, feature requests, and raising potential issues, while GitHub Discussions facilitate project-related discussions and knowledge sharing. Both the two sources on GitHub can provide architecture decisions from practical projects. Therefore, the three data sources ensure, to a certain extent, the diversity of our dataset. However, we acknowledge that our dataset is not comprehensive and does not cover all architecture problems. Another threat to the external validity is the selection of LLMs. We chose five widely used open-source and closed-source LLMs that have demonstrated strong performance in other SE research. However, the performance of these five LLMs may not fully represent the state-of-the-art capabilities of LLMs in generating DR, as other models with potentially superior performance may not be included. Therefore, due to the ever-evolving nature of LLMs, we plan to continue exploring their potential for generating DR in the future.
Reliability. Since the dataset construction and results evaluation process were conducted manually, there are potential to introduce personal bias. To reduce this threat, the first and fourth authors conducted pilot experiments before data labeling, extraction, and IHUM-category classification. During the data labeling and evaluation of the results (i.e., using accuracy metrics and IHUM-category), the consistency between the first and fourth authors was evaluated using the Cohen’s Kappa coefficient, yielding values of 0.834, 0.863, and 0.738, respectively. The Cohen’s Kappa value indicates an almost perfect agreement between the two authors on data labeling and accuracy metrics, and substantial agreement on the IHUM-category classification results (Landis and Koch, 1977). If any disagreement arose during the pilot and formal process, the second and third author were involved in reaching a consensus. The constructed dataset and evaluation results were reviewed multiple times by the four authors (the first to fourth authors) to ensure that they aligned with the pre-established criteria. Furthermore, we have made available the dataset of the study (Zhou et al., 2025) to enable other researchers to replicate and validate our findings.
7. Related Work
7.1. Design Rationale in Software Architecture
Several studies focused on DR in software architecture. Van der Ven et al. (van der Ven et al., 2006) indicated that explicitly modeling design decisions in software architecture can bridge the gap between rationale management and architecture artifacts, as it enables a close integration of rationale management with the architecture. Tang et al. (Tang et al., 2006) conducted a survey targeting practitioners to explore their perceptions of the value of DR and to understand how they utilize and document the background knowledge associated with their architecture decisions. In another study, Tang et al. (Tang et al., 2007) introduced a rationale-based architecture model that integrates design rationale, design objects, and their relationships to address the issue of design rationale often being undocumented or unstructured. Davide et al. (Falessi et al., 2008) proposed a new approach to Design Decision Rationale Documentation (DDRD), which tailors the documentation based on its intended use or purpose. In another study, Davide et al. (Falessi et al., 2013) conducted an empirical study through two controlled experiments and proved that the value of a Design Rationale Documentation (DRD) information item depends on its category (e.g., Assumptions, Related Requirements) and the activity it supports. Soliman et al. (Soliman, 2024) analyzed 156 architectural emails to identify the design rationale used in mailing lists and their relationship with decision types. They identified nine types of design rationale, six relationships between rationale types, and three relationships between decision and rationale types. Zhao et al. (Zhao et al., 2024) developed DRMiner, to automatically mine latent design rational from the discussion of developers in open-source community.
7.2. LLM-based Understanding of Software Artifacts
Numerous studies have been conducted to explore the capability of LLMs to understand various software artifacts. Xie et al. (Xie et al., 2024) evaluated the ability of LLMs to generate software specifications from software comments or documentation using few-shot learning. They compared the performance of 13 state-of-the-art LLMs with traditional methods across three publicly available datasets. Additionally, they conducted a comparative analysis of the failure cases from LLMs and traditional approaches, highlighting their respective strengths and weaknesses. Nam et al. (Nam et al., 2024) developed an LLM-based conversational user interface to assist users in understanding code. The results show that using their plugin is more effective in code completion than relying on web searches. Shaike et al. (Shaik et al., 2024) presented S3LLM, a framework based on LLMs, to understand large-scale scientific software using source code , code metadata, and summarized information from textual technical reports. Pex (Pex, 2024) studied the application of LLMs in software explanation and proposed the FLASE tool, which combines RAG and knowledge graphs, to enhance developers’ understanding of software systems. Pan et al. (Pan et al., 2025) conducted an empirical study to evaluate the capability of LLMs to understand design patterns. Franciscatto Guerra et al. (Guerra and Ernst, 2025) investigated the ability of an LLM to comprehend, replicate, and create structures within the intricate VIPER architecture, a design pattern used in iOS application development. Their results emphasize the potential of LLMs to lower development costs, as well as the challenges the LLMs face in being effectively applied to real-world software design situations. Casillo et al. (Casillo et al., 2025) constructed a dataset of 45,945 commits, each accompanied by rationale explaining the code changes. They then trained a model for generating rationale of code changes, highlighting the challenges in automating this process. Soliman et al. (Soliman and Keim, 2025) evaluated LLM’s understanding of software architecture by comparing its responses to a predefined ground truth. They found that while GPT offers initial insights, expert validation is still needed for reliable results of architecture knowledge. In their study, LLMs were also employed to generate DR for architecture decisions in HDFS.
7.3. Conclusive Summary
Previous studies (van der Ven et al., 2006; Tang et al., 2006, 2007) have highlighted the significant value of DR in software architecture. Besides, the application of LLM-based methods for understanding software artifacts (Xie et al., 2024; Pex, 2024; Soliman and Keim, 2025) has demonstrated the potential of LLMs in comprehending software architecture and generating DR. However, there has been limited focus on examining the capability of LLMs to generate DR for architectural decisions. Although the study of Soliman et al. (Soliman and Keim, 2025) covered querying LLMs about DR related to software architecture of HDFS, they mainly focused on examining whether LLMs possess specific architectural knowledge using the zero-shot method, rather than on the ability of LLMs to recover DR according to architectural problems and decisions. In our research, we placed greater emphasis on examining the capability of LLMs to generate DR for diverse real-world architecture problems and corresponding architecture decisions by collecting data from three sources (i.e., SO, GitHub Discussions, and GitHub Issues). Furthermore, we employed two complementary methods (accuracy metrics and the IHUM-category) to evaluate how the optimization of prompting strategies can enhance the quality of the generated DR by employing three different prompting strategies (i.e., zero-shot, CoT, and LLM-based agents).
8. Conclusions and Future Work
In this study, we analyzed the performance of LLMs in generating and recovering DR for architecture decisions. Our dataset consists of 50 SO posts, 25 GitHub issues, and 25 GitHub discussions that contain architecture decisions and corresponding DR. We used three prompting strategies (i.e., zero-shot, CoT, and LLM-based agent system) to instruct five selected LLMs to generate DR according to provided architecture problems and decisions. Finally, we evaluated the LLM-generated DR by both quantitative accuracy metrics (Precision, Recall, and F1-score) and qualitative analysis (IHUM-category). Our results show that, with the DR provided by human experts as the ground truth, the Precision of LLM-generated DR across the three prompting strategies ranges from 0.267 to 0.278, the Recall spans from 0.627 to 0.715, and the F1-score varies between 0.351 and 0.389. In addition, 64.45% to 69.42% of the arguments of generated DR not mentioned by human experts are also helpful, 4.12% to 4.87% of the arguments have uncertain correctness, and 1.59% to 3.24% of the arguments are potentially misleading. Compared to the zero-shot method, the CoT method generates DR with more arguments and explicitly considers trade-offs by weighing both advantages and disadvantages. However, the DR generated by the CoT method may lack in-depth analysis for individual arguments. The LLM-based agent system for DR generation acquires capabilities in information gathering, two layers of reasoning (i.e., Aspect_Reviewer and Trade-off_Analyst agents), and verification through multi-agent collaboration. Compared to the zero-shot and CoT methods, the DR generated by LLM-based agents is more aligned with the perspectives of human experts, and also contains fewer “Misleading” arguments. Additionally, the five LLMs can identify most of the key factors in the DR considered by human experts with the three prompting strategies and provide more comprehensive arguments. However, the five LLMs often struggle to provide analyses grounded in practical project experience and context-specific details as human experts do, and tend to include some “Uncertain” arguments, which are challenging to mitigate with the three employed prompting strategies.
With the evolution of LLMs, we plan to continuously explore the performance of emerging LLMs in DR generation. Additionally, according to the limitations observed in the five selected LLMs for DR generation, we aim to further optimize our LLM-based agent system. To be specific, we intend to enhance Information_Collector agent with the capability to locate code contexts relevant to architecture decisions within software projects, thereby providing Aspect_Analyst agent with specific code contexts for analysis. We also aim to augment Aspect_Analyst agent with the ability to identify the external knowledge underlying the generated arguments, as well as automatically generate code examples based on the arguments, to help LLM-based agents provide DR closer to practical development scenarios.
Data Availability
The replication package for this work has been made available at (Zhou et al., 2025).
Acknowledgements.
This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311 and 62402348.